Can someone help me here?
I need to solve the system by addition:
2x + 4y = 3
4x - 2y = 1
Thank you so much,
Lisa

Answer by pvaka(23)

2x + 4y = 3
4x - 2y = 1
2x + 4y = 3
2(4x + 2y) = (1) 2
2x + 4y = 3
8x - 4y = 2
=10x = 5
=10x/10 = 5/10
x=1/2
2(1/2) + 4y = 3
1 + 4y -1 = 3 - 1
4y = 2
4y/4 = 2/4
y=1/2
x = 1/2, y = 1/2
Hope this will help you out!